OpenStudy (anonymous):

The average height of an apple tree is 4.00 meters. How long would it take an apple falling from that height to reach the ground? Given: g = -9.8 meters/second^2

OpenStudy (anonymous):

Vf^2-Vi^2=2as Find Vf. Then use formula Vf=Vi+at to find t

OpenStudy (ajprincess):

use \(h=ut+\large\frac{1}{2}at^2\)

OpenStudy (anonymous):

t is?

OpenStudy (ajprincess):

t-time taken. that's vat u r askd to find.

OpenStudy (anonymous):

4 = 0 + 1/2 * 9.81 * t^2 8/9.81 = t^2 t =0.903

OpenStudy (anonymous):

The real question is whether anyone will hear the apple falling!!

OpenStudy (anonymous):

@Diarmuid_Wrenne WHAT!!

OpenStudy (anonymous):

@ajprincess that didnt work

OpenStudy (ajprincess):

vat did u get?

OpenStudy (anonymous):

is 4 the t?

OpenStudy (dls):

what's the answer anyway?

OpenStudy (anonymous):

@DLS thats not a answer choice

OpenStudy (dls):

Please do tell us the choices

OpenStudy (ajprincess):

no.  Nw solve to get t.